OpenStudy (anonymous):

Electronic baseball games manufactured by Tempco Electronics are shipped in lots of 24. Before shipping, a quality-control inspector randomly selects a sample of 8 from each lot for testing. If the sample contains any defective games, the entire lot is rejected. What is the probability that a lot containing exactly 3 defective games will still be shipped?

OpenStudy (anonymous):

the probability that the first one chosen is not defective is \(\frac{21}{24}\)

OpenStudy (anonymous):

the probability that the second one is not defective given that the first one is not, is \(\frac{20}{23}\)

OpenStudy (anonymous):

and finally the probability that the third one is not defective, given that the first two are not, is \(\frac{19}{22}\)

OpenStudy (anonymous):

multiply these together for your answer

OpenStudy (anonymous):

1.126???

OpenStudy (anonymous):

rounded three decimals places

OpenStudy (anonymous):

oh shoot did that wrong

OpenStudy (anonymous):

0.657

OpenStudy (anonymous):

?

OpenStudy (anonymous):

This seems wrong. Eight tests are run, and some of the later ones might find a defective.